OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10?

OpenStudy (jojo4eva):

slope=1/2

OpenStudy (jojo4eva):

what you would do is subtract 2x from both sides and get -4y=-2x+10 and then divide all terms by -4 and you would get y=-2/-4x-10/4 which equals y=1/2x-5/2

OpenStudy (anonymous):

Oh okay, thank you.